Start-up Observing Night 1 w/LBC 20160905 UT

Observers: JHill (LBTO), OKuhn(Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support: DMiller
Telescope Support: KNewton(LBTO)
Software Support: KSummers (Tucson)
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby

Summary

First night of startup observing using LBCs. Summary from Olga:

Last night (20160905UT) we obtained data to check the rotator center, dithering and guiding, zero points (skies were clear), and two superfoc sequences, one at PA=0 and the other at PA=180, to check focal plane tilt and focus offsets at the reference filters, V and r.

I haven't analyzed the data in detail, but examined the images, and everything looks fine - the first copoint (albeit before collimating) required only few-arcsec shifts and the series of superfoc sequences reached best-focus near the middle.

Although initial collimation was OK and converged after about 6 iterations, we ran into problems on the next field where the red pupils showed no inner hole and z22 was consistently negative ("the z22 run away"). To correct this, we tried WRS, but had some problems with it, and eventually, manual corrections helped. Marco was informed and is looking into the problems John reported.

LBC checks out OK. Will try to examine the data as soon as possible.

It would be nice to test the fix to IT 5715 (wait for offsets) once another instrument is available. The problem was that, when running pseudo-monocular with the other instrument dominant and LBC dithering, LBC images were started before the LBC offset had completed. LBC did not wait for the offset to complete when it did not send a preset (as it does not in pseudo-mono mode). Kellee fixed this in current version, v2.13.

Preparations

Issues with the interuption of LUCI2 precooling lead to a delay in the mounting of LBCs. And we had typical issues swapping LBC fibers.

Details

Open

04:40 Open to a beautiful clear sky which isn't quite photometric. Wind is 10-15 m/sec from South.

05:xx SA113

05:36 ACT0202 pointing star for lbcrangebal. Star was very close. Out of focus, mostly on blue.
       -------   -------   -------
   CE  -37.608     1.978   -35.630
   CA -109.506     1.477  -108.029
On LBC-Blue/LBC-Red the star is 8.6/9.6 arcsec away from the rotator center.Based on these, the following mount and optics offsets were calculated to co-point while optimizing the range available to M1: 
MOUNT: delta_IE = 1.98 arcsec, delta_CA = 1.48 arcsec
M1 SX: dX = 0.30 mm, dY = -0.24 mm, dRX = -5.17 arcsec, dRY = -6.47 arcsec 
M1 DX: dX = -0.30 mm, dY = 0.29 mm, dRX = 6.23 arcsec, dRY = 6.38 arcsec 

05:49 Slew to SA113, stop to run dofpia. _dofpia, /first_
Blue needs 21381 Z4 to start with, Red needs only -3500 Z4 at the start.
Converged after 6 iterations.

06:00: RotCenter Check:
SA113 with preset at PA=0 (060259b/060250r);
SA113 no preset PA=180 (060532b/060524r) and more (copying in the PA=0 OB and edited the dec, but my dither pattern was changed to 10 dithers of 0,0! This was the old OBGUI)
SA113 with preset PA=90 (061142b/061138r);
SA113 no preset PA=270 (061338b/061334r)
Red RC: 060250 & 060524:
      Old rotator center: 1078. 2913.
      New rotator center:
      ##################################################
      Xc=591.36 Yc=1467.78 from geomap.
      Xc=1080.69 Yc=2904.91 from averages.
      ##################################################
      Images are 2304 pixels wide

Red RC: 061138 & 061334:
      Old rotator center: 1078. 2913.
      New rotator center:
      ##################################################
      Xc=554.09 Yc=1399.35 from geomap.
      Xc=1080.29 Yc=2904.27 from averages.
      ##################################################
      Images are 2304 pixels wide

Blue RC: 060259 & 060532:
      Old rotator center: 1035. 2924.
      New rotator center:
      ##################################################
      Xc=569.15 Yc=1477.81 from geomap.
      Xc=1029.19 Yc=2919.48 from averages.
      ##################################################
      Images are 2304 pixels wide

Blue RC: 061142 & 061338:
      Old rotator center: 1035. 2924.
      New rotator center:
      ##################################################
      Xc=531.83 Yc=1407.64 from geomap.
      Xc=1030.4 Yc=2922.06 from averages.
      ##################################################
      Images are 2304 pixels wide

Rotator Centers (from averages, I did not investigate why geomap is giving such crazy results):
Channel RCx RCy OldRCx (from header) OldRCy (from header)
Red 1080.5 2904.5 1078 2913
Blue 1030 2921 1035 2924

At focal plane,13.5 mic/pixel, so the largest shift, 8.5 pix, between the new center and the value recorded in the header (the header is not changed for few-pixel shifts) corresponds to 0.114 mm.

06:16 Red channel IQ has gotten poor. dofpia again. Converged in 2 iterations, initial z4 corrections were ~2000-3000 nm.

06:21 Very clear ("1 level below really photometric" John says). SA113 photometric OB. AM = 1.187
  • banding on the red (62239). pickup noise on blue perhaps - very low level. Haven't ruled out aliasing in display.
  • DIMM bouncing around 1.3-1.5". FWHM of stars on chip2 is consistent with this. FWHM seem worse on edges, as usual, but we should run superfocs when seeing stabilizes.
  • Towards end of OB, after >~30 min, blue images are quite elongated.
  • Left/Right Preset cancelled (not by me or LBC) at 06:55...but the telescope is still tracing. This happened just before the last 1 or 2 images on each channel were taken.

06:58 Slewing to NGC7275 to run a "science" type OB which dither, switch filters, and guide on both channels.
  • Run NGC7252_CoPoint.ob.
  • dofpia. blue focus was apparently pretty close, but red was off by 4000nm. Both sides sent z22 = -100. Red is not converging --- it is repeatedly sending -100 z22. DIMM seeing (raw) ~ 1.4".
  • "S" to stop dofpia. John will run wrs to see whether it will help.

Corrected Red Z22 by hand.

07:05 Running DOWRS in "fast" mode (only correcting up to Z6). (still down at EL=32) Red asks me to click on pupil, but Blue did not. In both cases, the model didn't resemble the pupil. Blue's automated coma estimator does seem to be working. Hung after 'S', had to ^C.

Change WRS_RED.par to INTERACTIVE=0.

07:20 Trying DOHYBRID

Don't know what this one means:
% Wrong number of tags defined for structure: WRS.

This was a simple typo (keyword is WRSMode):
Reading the LBCFPIA config file
% Keyword WRS_MODE not allowed in call to: LBCFPIA_READ_CFG
% Execution halted at: DOHYBRID          620 /home/lbceng/WRS/src/dohybrid.pro
%                      $MAIN$          

07:32 Olga: We are slew to M31 copoint. dofpia. Shift "S" to stop so we can clear Active Optics and John can try dohybrid again (with typo corrected).

07:37 DOHYBRID (as lbceng) It successfully runs one pass of WRS, then switches to LBCFPIA. Then we seem to have a path problem.....

% Syntax error.
  At: /home/lbceng/WRS/src/anablob.pro, Line 36
% Compiled module: ANABLOB.
% Attempt to call undefined procedure/function: 'ANABLOB'.
% Execution halted at: FINDPUPILS        214 /home/lbceng/WRS/src/findpupils.pro
%                      LBCFPIA           243 /home/lbceng/WRS/src/lbcfpia.pro
%                      DOHYBRID          778 /home/lbceng/WRS/src/dohybrid.pro
%                      $MAIN$          
This was because /home/lbceng/WRS/lib/mpfit/ had all the files zipped up. I have unzipped them. But will that disturb WRS?

07:55: Back to dofpia on M31 copoint field. There is still so much spherical. Red is not converging - in a runaway negative z22. John tried to set gain=0 but that didn't take effect. Clear Active Optics again (08:07) and run dofpia again (08:08). John did turn off gain on z22. Red pupils still have spherical. Both red and blue are sending -100 z22, and neither is converging. "S" to test guiding but this is the dofpia problem that we need to work on.

08:20 Run copoint OB. Red IQ looks very bad (082106r). Copointing also required a bigger shift, probably due to all the mirror adjustments with "collimation".
          Old      Delta      New
       -------   -------   -------
   CE  -35.630   -16.418   -52.048
   CA -108.029    -3.757  -111.786
On LBC-Blue/LBC-Red the star is 16.5/36.7 arcsec away from the rotator center.Based on these, the following mount and optics offsets were calculated to co-point while optimizing the range available to M1: 
MOUNT: delta_IE = -16.42 arcsec, delta_CA = -3.76 arcsec
M1 SX: dX = 0.92 mm, dY = 0.57 mm, dRX = 12.24 arcsec, dRY = -19.69 arcsec 
M1 DX: dX = -0.91 mm, dY = -0.56 mm, dRX = -12.03 arcsec, dRY = 19.42 arcsec 

08:34. Despite poor collimation, running the OB to check that guiding works. The blue guide star is a donut, the red guide star looks like ghost-like. But guiding works. Stopped the OB... the IQ was just too poor for any 'pretty pictures'. We need to collimate & check focus offsets/tilt.

08:51 Slew to SA92. Copointing again since previous copoint made larger than expected adjustments. Focus/collimate on copoint field. Active optics were cleared prior to this, but z22 on red was turned off. On blue, it looked like negative z11 and z22 were needed, but on red, the pupil looked as before, with no inner hole discernible and it looks like positive z11 and z22 are needed. Stopping to try WRS on this.

08:55 DOWRS (as lbceng) FAST=0 This really doesn't do anything useful.

Changing to FAST=1

09:05 DOWRS (as lbceng) FAST=1

% Compiled module: ARRAY_INDICES.
% Compiled module: ZERNIKE_ESTIM.
***************** Zernike coeffs in mu starting from defocus **********************
      13.4436     0.328626    -0.742260   -0.0449259    0.0972854
***********************************************************************************
% Array subscript for ZE_TEMP must have same size as source expression.
% Execution halted at: CALL_WRS           98 /home/lbceng/WRS/src/call_wrs.pro
%                      LBCFPIA_WRS       256 /home/lbceng/WRS/src/lbcfpia_wrs.pro
%                      DOWRS             456 /home/lbceng/WRS/src/dowrs.pro
%                      $MAIN$          
% Program caused arithmetic error: Floating illegal operand

Clear Active Optics on both sides.

Change to FAST=0 SAVE=1

09:10 DOWRS

Preset is cancelled for telescope off source. El Axis was railed. Perhaps from wind+balance. Change fields back to north to get less wind.

09:20 DOWRS again

It looks like DOWRS is reading the wrong configuration file. It should be using lbcwrs.cfg.
% Compiled module: LBCFPIA_READ_CFG.
       1cfg file = ./lbcfpia.cfg
JMH fixed this the following day for DOHYBRID and DOWRS using an explicit /WRSMode in the call to LBCFPIA_READ_CFG.

09:34 "S" Stop DOWRS as the correction isn't getting any where,

10:06. With manual corrections and z22 gain off, we got pupils that looked good. But z22 was out of the threshold (50), so I used "S" to stop dofpia.

10:08 Without copointing (not to adjust optics any further), run through superfoc:
  • Vr PA=0
  • Vr PA=180

Results (Fri 9-9-2016). Best-focus on-axis:
SF_rfoc0      -0.0027 0.0006 18
SF_rfoc180  -0.0159 0.0037 20
SF_bfoc0     -0.0497 0.0058 75
SF_bfoc180 -0.1108 0.0199 32

and overall tilt:
rfoc0:      z =  2.300e-04 * x + -1.688e-04 * y + 5.030e+00
rfoc180:  z = -1.253e-04 * x + -6.022e-04 * y + 7.642e+00
bfoc0:     z = -2.878e-04 * x + -6.733e-04 * y + -6.528e+00
bfoc180: z = -1.707e-04 * x + 1.088e-04 * y + -7.055e+00

10:22 M31 copoint. 102325b/102321r
          Old      Delta      New
       -------   -------   -------
   CE  -52.048    21.226   -30.822
   CA -111.786    -1.436  -113.222
On LBC-Blue/LBC-Red the star is 9.1/36.3 arcsec away from the rotator center.Based on these, the following mount and optics offsets were calculated to co-point while optimizing the range available to M1: 
MOUNT: delta_IE = 21.23 arcsec, delta_CA = -1.44 arcsec
M1 SX: dX = -0.24 mm, dY = -0.70 mm, dRX = -14.99 arcsec, dRY = 5.17 arcsec 
M1 DX: dX = 0.25 mm, dY = 0.69 mm, dRX = 14.86 arcsec, dRY = -5.45 arcsec 

Do you want to send these offsets? Answer 'y' or 'Y' if so. Any other key will abort.

10:27 dofpia -- the copointing images had poor IQ. "S" because it is close, but not quite converging. Test guiding again, while on this M31 field and IQ is this time better than before.

10:45 Slew to SA95. dofpia. Converged in 2 iterations. Observing SA95. Airmass 1.267.

11:25 (18 deg twilight is at 11:33) Slew to Galadi field. dofpia. Red converged in 2, Blue in 3 iterations (except for Z22, whose gain is off --- "S" to stop dofpia). DIMM seeing is greater than 1" and highly variable between 1.5-1.

11:35 Start observing Galadi field. (OB is centered too S - the southern-most star is off the array).

11:45 Send updated OB with dec changed from 30 29 to 30 26. Running it from 2nd dither on. Still low. Adjust coordinates again before 3rd dither. 3rd dither started at 11:54, sky has brightened by about 1 mag. 00:29:21, 30:24:30 (J2000) center the cluster of stars well on chips 2,1 and 3. Offset to chip 4 should be made less than 772, but only 1-2 stars off top of chip 4.

Sky Flats

12:09 Slew to SkyFlats position (HA=00:50, DEC=27). Sun @ -10deg. Not bright enough yet.

12:24 -7.4deg. 4-sec VV. PA=0 (31k...58k for Blue and 21k...41k for Red).

12:30 -6.4deg 1-sec Br PA=0 (20k...61k for R)

12:34 0.5-sec Bz PA=0. (4 of 5 dithers only)

12:37 Sun @ -4.9deg. 0.7...0.5-sec SY PA=0 (the set of Y flats is OK, 26-42k counts, but the set of S runs from 40-56k)

Close

12:41 Slew to horizon to close shutter doors.

-- %USERSIG{JohnHill - 2016-09-05}%
Topic revision: r14 - 09 Sep 2016, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback